# Constructions of binary self-orthogonal singly-even minimal linear codes violating the Aschikhmin-Barg condition with few weights

### v2

@@ 5 Constructions of binary SO singly-even minimal linear codes from vectorial Boolean functions
TABLE III:  The values of $W_{F}(v,u,b_{1},b_{2})$ when $t$ is even and $(\lambda_{1}+\lambda_{2})^{-1}=\lambda_{1}^{-1}+\lambda_{2}^{-1}$
Value
Condition
$2^{m}$ $v=u=b_{1}=b_{2}=0$
0 $v=u=0,(b_{1},b_{2})\in(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*}$
$(-1)^{\mathrm{i}}2^{t}-4$ $v=0,u=1,(b_{1},b_{2})\in\mathcal{V}\times\mathbb{F}_{2^{t}}$, $\mathcal{A}_{3}=\mathrm{i}$
$3\cdot 2^{t-1}-2^{m-1}-4$ $v=u=1,b_{1}=b_{2}=0$
$3\cdot 2^{t-1}-4$ $v=u=1,(b_{1},b_{2})\in(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*},(\mathcal{A}_{1},\mathcal{A}_{2})=(0,0)$
$2^{t-1}-4$ $v=u=1,(b_{1},b_{2})\in(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*},(\mathcal{A}_{1},\mathcal{A}_{2})\in\{(\mathrm{i}+1,\mathrm{i}),(1,1)\}$
$2^{m-1}+2^{t-1}$ $v=1,u=0,(b_{1},b_{2})=(0,0)$
$3\cdot 2^{t-1}$ $v=1,u=0,(b_{1},b_{2})\in(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*},(\mathcal{A}_{1},\mathcal{A}_{2})=(1,1)$
$2^{t-1}$ $v=1,u=0,(b_{1},b_{2})\in(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*},(\mathcal{A}_{1},\mathcal{A}_{2})\in\{(\mathrm{i}+1,\mathrm{i}),(0,0)\}$
TABLE IV:  The values of $W_{F}(v,u,b_{1},b_{2})$ when $t$ is odd, or $t$ is even and $(\lambda_{1}+\lambda_{2})^{-1}\neq\lambda_{1}^{-1}+\lambda_{2}^{-1}$
Value
Condition
$2^{m}$ $v=u=b_{1}=b_{2}=0$
0 $v=u=0,(b_{1},b_{2})\in(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*}$
$(-1)^{\mathrm{i}}2^{t}-4$ $v=0,u=1,(b_{1},b_{2})\in\mathcal{V}\times\mathbb{F}_{2^{t}}$, $\mathcal{A}_{3}=\mathrm{i}$
$3\cdot 2^{t-1}-2^{m-1}-4$ $v=u=1,b_{1}=b_{2}=0$
$(-1)^{\mathrm{i}}3\cdot 2^{t-1}-4$ $v=u=1,(b_{1},b_{2})\in(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*},{\rm wt}(\mathcal{A}_{1},\mathcal{A}_{2},\mathcal{A}_{3})=3\mathrm{i}$
$(-1)^{\mathrm{i}}2^{t-1}-4$ $v=u=1,(b_{1},b_{2})\in(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*},{\rm wt}(\mathcal{A}_{1},\mathcal{A}_{2},\mathcal{A}_{3})=1+\mathrm{i}$
$2^{m-1}+2^{t-1}$ $v=1,u=0,(b_{1},b_{2})=(0,0)$
$(-1)^{\mathrm{i}}2^{t-1}$
$v=1,u=0,(b_{1},b_{2})\in(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*},(\mathcal{A}_{1},\mathcal{A}_{2},\mathcal{A}_{3})\in\{(\mathrm{i},\mathrm{i},\mathrm{i}),(1+\mathrm{i},\mathrm{i},1+\mathrm{i}),(\mathrm{i},1+\mathrm{i},1+\mathrm{i})\}$
$(-1)^{\mathrm{i}}3\cdot 2^{t-1}$ $v=1,u=0,(b_{1},b_{2})\in(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*},(\mathcal{A}_{1},\mathcal{A}_{2},\mathcal{A}_{3})=(\mathrm{i},\mathrm{i},1+\mathrm{i})$
@@ 0 $v=u=0,(b_{1},b_{2})\in(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*}$
For any given $(\mu=(v,u),b_{1},b_{2})\in\mathbb{F}_{2}\times\mathbb{F}_{2}\times\mathcal{V}\times\mathbb{F}_{2^{t}}$, we have
$W_{F}(\mu,b_{1},b_{2})$ $=$ $\sum\limits_{(y,x)\in\mathbb{F}_{2^{t}}\times\mathbb{F}_{2^{t}}}(-1)^{vf(y,x)+uG(y,x)+\mathrm{Tr}_{2^{t}}(b_{1}x)+\mathrm{Tr}_{2^{t}}(b_{2}y)}$ (38)
$=$ $\left\{\begin{array}[]{ll}W_{G}(u,b_{1},b_{2}),&\mathrm{if}~v=0,\\ W_{H_{u}}(b_{1},b_{2}),&\mathrm{if}~v=1.\end{array}\right$.
We first calculate the value of $W_{G}(u,b_{1},b_{2})$ for any $(u,b_{1},b_{2})\in\mathbb{F}_{2}\times\mathcal{V}\times\mathbb{F}_{2^{m}}$. By Eq. (38), we obtain
$W_{G}(u,b_{1},b_{2})$ (43)
$=$ $\sum\limits_{(y,x)\in\mathbb{F}_{2^{t}}\times\mathbb{F}_{2^{t}}}(-1)^{u(\mathrm{Tr}_{2^{t}}(\lambda_{3}y\pi(x))+((x+w_{1})^{2^{m}-1}+(x+w_{2})^{2^{m}-1})(y^{2^{m}-1}+1))+\mathrm{Tr}_{2^{t}}(b_{1}x+b_{2}y)}$
$=$ $\sum\limits_{(y,x)\in\mathbb{F}_{2^{t}}\times\mathbb{F}_{2^{t}}\setminus\{(0,w_{1}),(0,w_{2})\}}(-1)^{u\mathrm{Tr}_{2^{t}}(\lambda_{3}y\pi(x))+\mathrm{Tr}_{2^{t}}(b_{1}x+b_{2}y)}+(-1)^{u}\sum\limits^{2}_{i=1}(-1)^{\mathrm{Tr}_{2^{t}}(b_{1}w_{i})}$
$=$ $\sum\limits_{(y,x)\in\mathbb{F}_{2^{t}}\times\mathbb{F}_{2^{t}}}(-1)^{u\mathrm{Tr}_{2^{t}}(\lambda_{3}y\pi(x))+\mathrm{Tr}_{2^{t}}(b_{1}x+b_{2}y)}-(1-(-1)^{u})\sum\limits^{2}_{i=1}(-1)^{\mathrm{Tr}_{2^{t}}(b_{1}w_{i})}$
$=$ $\left\{\begin{array}[]{ll}2^{m},&\mathrm{if}~u=b_{1}=b_{2}=0,\\ 0,&\mathrm{if}~u=0,~(b_{1},b_{2})\in(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*},\\ 2^{t}-4,&\mathrm{if}~u=1,b_{1}=b_{2}=0,\\ (-1)^{\mathrm{i}}2^{t}-4,&\mathrm{if}~u=1,(b_{1},b_{2})\in(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*},A_{3}=\mathrm{i}.\end{array}\right$.
Note that if $t$ is odd, then Lemma 5.6 implies that $(\lambda_{1}+\lambda_{2})^{-1}\neq\lambda_{1}^{-1}+\lambda_{2}^{-1}$. Furthermore, we shall divide the computation of the values of $W_{H_{1}}(b_{1},b_{2})$ into two cases according to whether $(\lambda_{1}+\lambda_{2})^{-1}$ is equal to $\lambda_{1}^{-1}+\lambda_{2}^{-1}$.
(1) If $t$ is even and $(\lambda_{1}+\lambda_{2})^{-1}=\lambda_{1}^{-1}+\lambda_{2}^{-1}$, then by the method analogous to that used in Eq. (26), we have
$W_{H_{1}}(b_{1},b_{2})$
$=$ $\left\{\begin{array}[]{ll}3\cdot 2^{t-1}-2^{m-1}-4,&\mathrm{if}~b_{1}=b_{2}=0,\\ 3\cdot 2^{t-1}-4,&\mathrm{if}~(b_{1},b_{2})\in(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*},(\mathcal{A}_{1},\mathcal{A}_{2})=(0,0),\\ -2^{t-1}-4,&\mathrm{if}~(b_{1},b_{2})\in(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*},(\mathcal{A}_{1},\mathcal{A}_{2})\in\{(\mathrm{i}+1,\mathrm{i}),(1,1)\}.\end{array}\right$.
(2) If $t$ is odd, or $t$ is even and $(\lambda_{1}+\lambda_{2})^{-1}\neq\lambda_{1}^{-1}+\lambda_{2}^{-1}$, then by the method analogous to that used in Eq. (31), we have
$W_{H_{1}}(b_{1},b_{2})$
$=$ $\left\{\begin{array}[]{ll}3\cdot 2^{t-1}-2^{m-1}-4,&\mathrm{if}~b_{1}=b_{2}=0,\\ (-1)^{\mathrm{i}}3\cdot 2^{t-1}-4,&\mathrm{if}~(b_{1},b_{2})\in(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*},{\rm wt}(\mathcal{A}_{1},\mathcal{A}_{2},\mathcal{A}_{3})=3\mathrm{i}\\ (-1)^{\mathrm{i}}2^{t-1}-4,&\mathrm{if}~(b_{1},b_{2})\in(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*},{\rm wt}(\mathcal{A}_{1},\mathcal{A}_{2},\mathcal{A}_{3})=1+\mathrm{i}.\end{array}\right$.
Finally, the values of $W_{f}(b_{1},b_{2})=W_{H_{0}}(b_{1},b_{2})$ follow directly from Eqs. (26) and (31). The desired result is thus obtained by combining the above arguments. ∎
Applying Lemma 5.8 to Lemma 5.2, we can deduce the following result.
Theorem 5.9.
Let $m=2t$, notation and conditions be the same as in Lemma 5.8. Let $\mathcal{C}_{F}$ be the linear code derived by
${\mathcal{C}}_{F}=\left\{{\bf u}_{a,b}=((v,u)\cdot F(y,x)+b_{1}y+b_{2}x)_{(y,x)\in{\mathbb{F}}_{2}^{t}\times{\mathbb{F}}_{2}^{t}\backslash\{(\mathbf{0},\mathbf{0})\}}:(v,u)\in{\mathbb{F}}_{2}\times{\mathbb{F}}_{2},(b_{1},b_{2})\in\mathcal{V}\times{\mathbb{F}}_{2^{t}}\right\},$
@@
  1. (1)
  2. (2)
@@
Plugging the values of $W_{F}(v,u,b_{1},b_{2})$ from Tables III and IV back into Lemma 2.3, we obtain the weights of $\mathcal{C}_{F}$ in Tables V and VI, respectively. Denote by $N$ the number of $(b_{1},b_{2})$ satisfying $W_{F}(v,u,b_{1},b_{2})=3\cdot 2^{m-1}-4$ in $(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*}$, we now proceed to calculate $N$. Incorporated with Definition 5.5 and Table III, it follows that
$N$
$=$ $\sum\limits_{b_{1}\in\mathrm{V}}\sum\limits_{b_{2}\in\mathbb{F}_{2^{t}}}\left(\frac{1+(-1)^{\mathcal{A}_{1}}}{2}\right)\left(\frac{1+(-1)^{\mathcal{A}_{2}}}{2}\right)-1$
$=$ $\sum\limits_{b_{1}\in\mathbb{F}_{2^{t}}}\sum\limits_{b_{2}\in\mathbb{F}_{2^{t}}}\left(\frac{1+(-1)^{\mathcal{A}_{1}}}{2}\right)\left(\frac{1+(-1)^{\mathcal{A}_{2}}}{2}\right)\prod^{2}_{i=1}\left(\frac{1+(-1)^{\mathrm{Tr}_{2^{t}}\left(b_{1}w_{i}\right)}}{2}\right)-1$
$=$ $2^{m-4}+3\cdot 2^{t-2}-1.$
The frequency of the other cases in Table III and all cases in Table VI can be obtained by the method analogous to that used above. By applying Tables III and IV to Lemmas 3.6 and 5.2, we can deduce that the codes $\mathcal{C}_{F}$ are SO singly-even, minimal, and violate the AB condition. ∎
@@
$(1)$ Let $m=2t=8$, define $F(y,x)=(\mathrm{Tr}_{2^{t}}(\xi yx^{2})\mathrm{Tr}_{2^{t}}(\xi^{6}yx^{2}),\mathrm{Tr}_{2^{t}}((\xi+\xi^{6})y\pi(x))+((x+\xi)^{2^{t}-1}+(x+\xi^{2})^{2^{t}-1})(y^{2^{t}-1}+1))$, where $\xi$ is a primitive element in $\mathbb{F}_{2^{4}}$ such that $\xi^{4}+\xi+1=0$. Then by MAGMA we have $(\xi+\xi^{6})^{-1}=\xi^{-1}+\xi^{-6}$. Let $\mathcal{V}=\{x\in\mathbb{F}_{2^{t}}:\mathrm{Tr}_{2^{t}}(\xi x)=\mathrm{Tr}_{2^{t}}(\xi^{2}x)=0\}$. Then we can verify by MAGMA that the linear code $\mathcal{C}_{F}$ is a binary SO singly-even minimal $[255,8,60]$-linear code violating the AB condition with the weight enumerator is $1+z^{60}+27z^{118}+40z^{122}+51z^{124}+63z^{128}+36z^{134}+24z^{138}+12z^{140}+z^{182}$, which is consistent with Theorem 5.9(1).
$(2)$ Let $m=2t=10$, and define $F(y,x)=(\mathrm{Tr}_{2^{t}}(\xi yx^{2})\mathrm{Tr}_{2^{t}}(\xi^{2}yx^{2}),\mathrm{Tr}_{2^{t}}((\xi+\xi^{2})y\pi(x))+((x+\xi)^{2^{t}-1}+(x+\xi^{2})^{2^{t}-1})(y^{2^{t}-1}+1))$, where $\xi$ is a primitive element in $\mathbb{F}_{2^{5}}$ such that $\xi^{5}+\xi^{2}+1=0$. Let $\mathcal{V}=\{x\in\mathbb{F}_{2^{t}}:\mathrm{Tr}_{2^{t}}(\xi x)=\mathrm{Tr}_{2^{t}}(\xi^{2}x)=0\}$. Then, we can verify by MAGMA that the linear code $\mathcal{C}_{F}$ is a binary SO singly-even minimal $[1023,10,248]$-linear code violating the AB condition with the weight enumerator is $1+z^{248}+28z^{488}+59z^{490}+144z^{498}+115z^{504}+84z^{506}+255z^{512}+84z^{520}+84z^{522}+112z^{530}+28z^{536}+28z^{538}+z^{746}$, which is consistent with Theorem 5.9(2).

### v3

@@ 5 Constructions of binary SO singly-even minimal linear codes from vectorial Boolean functions
TABLE III:  The values of $W_{F}(a_{1},a_{2},b_{1},b_{2})$ when $(\lambda_{1}+\lambda_{2})^{-1}=\lambda_{1}^{-1}+\lambda_{2}^{-1}$
Value Condition
$2^{m}$ $a_{1}=a_{2}=b_{1}=b_{2}=0$
0 $a_{1}=a_{2}=0,(b_{1},b_{2})\in(\mathbb{F}_{2^{t}}\times\mathcal{V})^{*}$
$(-1)^{\mathrm{i}}2^{t}-4$ $a_{1}=0,a_{2}=1,(b_{1},b_{2})\in\mathbb{F}_{2^{t}}\times\mathcal{V}$, $\mathcal{A}_{3}=\mathrm{i}$
$3\cdot 2^{t-1}-2^{m-1}-4$ $a_{1}=a_{2}=1,b_{1}=b_{2}=0$
$3\cdot 2^{t-1}-4$ $a_{1}=a_{2}=1,(b_{1},b_{2})\in(\mathbb{F}_{2^{t}}\times\mathcal{V})^{*},(\mathcal{A}_{1},\mathcal{A}_{2})=(0,0)$
$2^{t-1}-4$ $a_{1}=a_{2}=1,(b_{1},b_{2})\in(\mathbb{F}_{2^{t}}\times\mathcal{V})^{*},(\mathcal{A}_{1},\mathcal{A}_{2})\in\{(\mathrm{i}+1,\mathrm{i}),(1,1)\}$
$2^{m-1}+2^{t-1}$ $a_{1}=1,a_{2}=0,(b_{1},b_{2})=(0,0)$
$3\cdot 2^{t-1}$ $a_{1}=1,a_{2}=0,(b_{1},b_{2})\in(\mathbb{F}_{2^{t}}\times\mathcal{V})^{*},(\mathcal{A}_{1},\mathcal{A}_{2})=(1,1)$
$2^{t-1}$ $a_{1}=1,a_{2}=0,(b_{1},b_{2})\in(\mathbb{F}_{2^{t}}\times\mathcal{V})^{*},(\mathcal{A}_{1},\mathcal{A}_{2})\in\{(\mathrm{i}+1,\mathrm{i}),(0,0)\}$
TABLE IV:  The values of $W_{F}(v,u,b_{1},b_{2})$ when $(\lambda_{1}+\lambda_{2})^{-1}\neq\lambda_{1}^{-1}+\lambda_{2}^{-1}$
Value Condition
$2^{m}$ $a_{1}=a_{2}=b_{1}=b_{2}=0$
0 $a_{1}=a_{2}=0,(b_{1},b_{2})\in(\mathbb{F}_{2^{t}}\times\mathcal{V})^{*}$
$(-1)^{\mathrm{i}}2^{t}-4$ $a_{1}=0,a_{2}=1,(b_{1},b_{2})\in\mathbb{F}_{2^{t}}\times\mathcal{V}$, $\mathcal{A}_{3}=\mathrm{i}$
$3\cdot 2^{t-1}-2^{m-1}-4$ $a_{1}=a_{2}=1,b_{1}=b_{2}=0$
$(-1)^{\mathrm{i}}3\cdot 2^{t-1}-4$ $a_{1}=a_{2}=1,(b_{1},b_{2})\in(\mathbb{F}_{2^{t}}\times\mathcal{V})^{*},{\rm wt}(\mathcal{A}_{1},\mathcal{A}_{2},\mathcal{A}_{3})=3\mathrm{i}$
$(-1)^{\mathrm{i}}2^{t-1}-4$ $a_{1}=a_{2}=1,(b_{1},b_{2})\in(\mathbb{F}_{2^{t}}\times\mathcal{V})^{*},{\rm wt}(\mathcal{A}_{1},\mathcal{A}_{2},\mathcal{A}_{3})=1+\mathrm{i}$
$2^{m-1}+2^{t-1}$ $a_{1}=1,a_{2}=0,(b_{1},b_{2})=(0,0)$
$(-1)^{\mathrm{i}}2^{t-1}$ $a_{1}=1,a_{2}=0,(b_{1},b_{2})\in(\mathbb{F}_{2^{t}}\times\mathcal{V})^{*},(\mathcal{A}_{1},\mathcal{A}_{2},\mathcal{A}_{3})\in\{(\mathrm{i},\mathrm{i},\mathrm{i}),(1+\mathrm{i},\mathrm{i},1+\mathrm{i}),(\mathrm{i},1+\mathrm{i},1+\mathrm{i})\}$
$(-1)^{\mathrm{i}}3\cdot 2^{t-1}$ $a_{1}=1,a_{2}=0,(b_{1},b_{2})\in(\mathbb{F}_{2^{t}}\times\mathcal{V})^{*},(\mathcal{A}_{1},\mathcal{A}_{2},\mathcal{A}_{3})=(\mathrm{i},\mathrm{i},1+\mathrm{i})$
@@ 0 $a_{1}=a_{2}=0,(b_{1},b_{2})\in(\mathbb{F}_{2^{t}}\times\mathcal{V})^{*}$
According to the definition of the Walsh transform, for any $(a_{1},a_{2},b_{1},b_{2})\in\mathbb{F}_{2}\times\mathbb{F}_{2}\times\mathbb{F}_{2^{t}}\times\mathcal{V}$, we have
$W_{F}(a_{1},a_{2},b_{1},b_{2})$ $=$ $\sum_{(x,y)\in{\mathbb{F}}_{2^{t}}^{2}}(-1)^{a_{1}f(x,y)+a_{2}g(x,y)+{\rm Tr}(b_{1}x+b_{2}y)}$ (29)
@@
Case I: If $\lambda_{3}^{-1}=\lambda_{1}^{-1}+\lambda_{2}^{-1}$, i.e., ${\mathcal{A}}_{3}={\mathcal{A}}_{1}+{\mathcal{A}}_{2}$, then by Lemma 5.6, we have
$W_{f}(b_{1},b_{2})=\begin{cases}2^{m-1}+2^{t-1},&~\text{if}~(b_{1},b_{2})=(0,0),\\ -3\cdot 2^{t-1},&~\text{if}~(b_{1},b_{2})\neq(0,0)~\text{and}~({\mathcal{A}}_{1},{\mathcal{A}}_{2})=(1,1),\\ -2^{t-1},&~\text{if}~(b_{1},b_{2})\neq(0,0)~\text{and}~({\mathcal{A}}_{1},{\mathcal{A}}_{2})\in\{(0,0),(1,0),(0,1)\}\end{cases}$
and
[…]
Then with Eq. (29), we can obtain all cases in Table III.
Case II: If $\lambda_{3}^{-1}\neq\lambda_{1}^{-1}+\lambda_{2}^{-1}$, i.e., ${\mathcal{A}}_{3}={\mathcal{A}}_{1}+{\mathcal{A}}_{2}$ does not always hold for any $(b_{1},b_{2})\in(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*}$, all cases in Table IV can be obtained by Eq. (29) and Lemma 5.6. ∎
Remark 5.9.
In Lemma 5.8, $(\lambda_{1}+\lambda_{2})^{-1}=\lambda_{1}^{-1}+\lambda_{2}^{-1}$ if and only if $\lambda_{1}^{2}+\lambda_{2}^{2}=\lambda_{1}\lambda_{2}$, which holds if and only if $t$ is even and $\frac{\lambda_{1}}{\lambda_{2}}\in{\mathbb{F}}_{4}\backslash{\mathbb{F}}_{2}$.
@@
Applying Lemmas 5.7 and 5.8 to Lemma 5.2, we can deduce the following result.
Theorem 5.10.
Let $m=2t\geq 6$, notation and conditions be the same as in Lemma 5.8. Let $\mathcal{C}_{F}$ be the linear code derived by
${\mathcal{C}}_{F}=\left\{((a_{1},a_{2})\cdot F(x,y)+b_{1}x+b_{2}y)_{(x,y)\in{\mathbb{F}}_{2}^{t}\times{\mathbb{F}}_{2}^{t}\backslash\{(\mathbf{0},\mathbf{0})\}}:(a_{1},a_{2})\in{\mathbb{F}}_{2}\times{\mathbb{F}}_{2},(b_{1},b_{2})\in{\mathbb{F}}_{2^{t}}\times{\mathcal{V}}\right\},$
where $F(x,y)$ is defined in Definition 5.5. Then $\mathcal{C}_{F}$ is a $2^{m}-1,m,2^{m-2}-2^{t-2}]-$linear code. Furthermore, the following two statements hold.
  1. (1)
  2. (2)
@@
Plugging the values of $W_{F}$ from Tables III and IV back into Lemma 2.6, we can obtain the weights of $\mathcal{C}_{F}$ in Tables V and VI, respectively. As for an example, we compute the number, denoted by $N$, of $(b_{1},b_{2})\in(\mathbb{F}_{2^{t}}\times\mathcal{V})^{*}$ satisfying $W_{F}(1,1,b_{1},b_{2})=3\cdot 2^{t-1}-4$ in $(\mathbb{F}_{2^{t}}\times\mathcal{V})^{*}$, i.e, $({\mathcal{A}}_{1},{\mathcal{A}}_{2})=(0,0)$ from Table III. By the basic knowledge of the exponential sum, we have
$N$ $=$ $\frac{1}{4}\sum_{b_{1}\in{\mathbb{F}}_{2^{t}}}\sum_{b_{2}\in{\mathcal{V}}}\sum_{s_{1}\in{\mathbb{F}}_{2}}(-1)^{{\mathcal{A}}_{1}s_{1}}\sum_{s_{2}\in{\mathbb{F}}_{2}}(-1)^{{\mathcal{A}}_{2}s_{2}}-1$
$=$ $\frac{1}{4}\sum_{b_{1}\in{\mathbb{F}}_{2^{t}}}\sum_{b_{2}\in{\mathcal{V}}}\left(1+(-1)^{{\mathcal{A}}_{1}}\right)\left(1+(-1)^{{\mathcal{A}}_{2}}\right)-1$
$=$ $\frac{1}{4}\sum_{b_{1}\in{\mathbb{F}}_{2^{t}}}\sum_{b_{2}\in{\mathcal{V}}}\left(1+(-1)^{{\mathcal{A}}_{1}}+(-1)^{{\mathcal{A}}_{2}}+(-1)^{{\mathcal{A}}_{3}}\right)-1$
$=$ $2^{m-4}+3\cdot 2^{t-2}-1,$
where the last equality holds due to the fact that for all $i\in\{1,2,3\},$
$\sum_{b_{1}\in{\mathbb{F}}_{2^{t}}}\sum_{b_{2}\in{\mathcal{V}}}(-1)^{{\mathcal{A}}_{i}}=\sum_{b_{2}\in{\mathcal{V}}}\sum_{b_{1}\in{\mathbb{F}}_{2^{t}}}(-1)^{{\rm Tr}(b_{2}\pi^{-1}(\lambda_{i}^{-1}b_{1}))}=2^{t}.$
The frequency of the other cases in Table III and all cases in Table VI can be obtained by the method analogous to that used above.
By applying Tables III and IV to Lemmas 3.6 and 5.2, we can deduce that the codes $\mathcal{C}_{F}$ are SO singly-even, minimal, and violate the AB condition. ∎
@@
$(1)$ Let $m=2t=8$, define $F(x,y)=(\mathrm{Tr}(\xi xy^{2})\mathrm{Tr}(\xi^{6}xy^{2}),\mathrm{Tr}((\xi+\xi^{6})x\pi(y))+((y+\xi)^{2^{t}-1}+(y+\xi^{2})^{2^{t}-1})(x^{2^{t}-1}+1))$, where $\xi$ is a primitive element in $\mathbb{F}_{2^{4}}$ such that $\xi^{4}+\xi+1=0$. Then by MAGMA, we have $(\xi+\xi^{6})^{-1}=\xi^{-1}+\xi^{-6}$. Let $\mathcal{V}=\{x\in\mathbb{F}_{2^{t}}:\mathrm{Tr}(\xi x)=\mathrm{Tr}(\xi^{2}x)=0\}$. Then we can verify by MAGMA that the linear code $\mathcal{C}_{F}$ is a binary SO singly-even minimal $[255,8,60]$-linear code violating the AB condition with the weight enumerator is $1+z^{60}+27z^{118}+40z^{122}+51z^{124}+63z^{128}+36z^{134}+24z^{138}+12z^{140}+z^{182}$, which is consistent with Theorem 5.10(1).
$(2)$ Let $m=2t=10$, and define $F(x,y)=(\mathrm{Tr}(\xi xy^{2})\mathrm{Tr}(\xi^{2}xy^{2}),\mathrm{Tr}((\xi+\xi^{2})x\pi(y))+((y+\xi)^{2^{t}-1}+(y+\xi^{2})^{2^{t}-1})(x^{2^{t}-1}+1))$, where $\xi$ is a primitive element in $\mathbb{F}_{2^{5}}$ such that $\xi^{5}+\xi^{2}+1=0$. Let $\mathcal{V}=\{x\in\mathbb{F}_{2^{t}}:\mathrm{Tr}(\xi x)=\mathrm{Tr}(\xi^{2}x)=0\}$. Then, we can verify by MAGMA that the linear code $\mathcal{C}_{F}$ is a binary SO singly-even minimal $[1023,10,248]$-linear code violating the AB condition with the weight enumerator is $1+z^{248}+28z^{488}+59z^{490}+144z^{498}+115z^{504}+84z^{506}+255z^{512}+84z^{520}+84z^{522}+112z^{530}+28z^{536}+28z^{538}+z^{746}$, which is consistent with Theorem 5.10(2).